# AST-T5: Structure-Aware Pretraining for Code Generation and Understanding

### v1

@@ Abstract
Large language models (LLMs) have made significant advancements in code-related tasks, yet many LLMs treat code as simple sequences, neglecting its structured nature. We introduce AST-T5, a novel pretraining paradigm that leverages the Abstract Syntax Tree (AST) for enhanced code generation, transpilation, and understanding. Using dynamic programming, our AST-Aware Segmentation retains code structure, while our AST-Aware Span Corruption objective equips the model to reconstruct various code structures. Unlike other models, AST-T5 avoids intricate program analyses or architectural changes, so it integrates seamlessly with any encoder-decoder Transformer. Evaluations show that AST-T5 consistently outperforms similar-sized LMs across various code-related tasks. Structure-awareness makes AST-T5 particularly powerful in code-to-code tasks, surpassing CodeT5 by 2 points in exact match score for the Bugs2Fix task and by 3 points in exact match score for Java-C# Transpilation in CodeXGLUE. Our code and model are publicly available at https://github.com/gonglinyuan/ast_t5.
@@ 1 Introduction
We have witnessed the transformative impact of large language models (LLMs) on various aspects of artificial intelligence in recent years (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023), especially in code generation and understanding (Feng et al., 2020; Wang et al., 2021; Rozière et al., 2023). By pretraining on massive code corpora such as the GitHub corpus, LLMs learns rich representations, thereby becoming powerful tools for various downstream applications, including text-to-code generation (Chen et al., 2021a; Austin et al., 2021; Iyer et al., 2018), code-to-code transpilation (Lu et al., 2021; Lachaux et al., 2020; Tufano et al., 2019), and code understanding (mapping code to classification labels) (Zhou et al., 2019; Svajlenko et al., 2014).
@@
Figure 1: Comparison of AST-Aware Subtree Corruption and Vanilla T5 using a Python factorial function. Both methods replace masked spans with sentinel tokens (special tokens added to the vocabulary, shown as [X], [Y], and [Z] in the figure), with output sequences containing the original masked tokens. Inputs and targets are shown in byte-pair encoding (BPE); for instance, “factorial” is encoded into “fact” and “##orial”. Unlike Vanilla T5, which masks random spans without considering code structure, our approach specifically targets spans aligned with AST subtrees, like expressions and statements.
In our experiments, AST-T5 consistently outperforms baselines in code generation, transpilation, and understanding tasks. Through controlled experiments, we empirically demonstrate that these advancements are attributed to our AST-aware pretraining techniques. Notably, AST-T5 not only outperforms similar-sized models like CodeT5 and CodeT5+ across various benchmarks but also remains competitive with, or occasionally even exceeds, the performance of much larger models using the HumanEval dataset. Furthermore, the inherent AST-awareness of AST-T5 offers unique advantages in structure-sensitive tasks, such as code-to-code transpilation and Clone Detection, highlighting its effectiveness at capturing the structural nuances of code.
@@ 3.1 Parsing Code Into ASTs
Unlike traditional language models on code that handle code as simple sequences of subword tokens, AST-T5 leverages the Abstract Syntax Tree (AST) of code to gain semantic insights. For parsing purposes, we assume the provided code is syntactically valid—a reasonable assumption for tasks like code transpilation and understanding. Instead of the often computationally-intensive or infeasible methods of Control-Flow Analysis (CFA) or code execution (Guo et al., 2021; Tipirneni et al., 2023), our method only demands the code to be parsable. We use Tree-sitter, a multi-language parser, to construct the ASTs, where each subtree represents a consecutive span of subword tokens, and every leaf node represents an individual token.
3.2 Our AST-Aware Segmentation
@@
Segmentation in language model pretraining
is a critical yet often overlooked aspect. Transformer LMs impose token limits on input sequences, making segmentation essential for fitting these inputs within the max_len constraint. A naive approach is Greedy Segmentation, where each chunk, except the last, contains exactly max_len tokens Figure 2 (Left). This strategy has been widely adopted in previous works, such as CodeT5 (Wang et al., 2021).
@@
AST-Aware Segmentation
is our novel approach designed to preserve the AST structure of code during segmentation. Unlike Greedy Segmentation, which can indiscriminately fragment AST structures, our method strategically minimizes such disruptions. As illustrated in the example in Figure 2, Greedy Segmentation leads to nine instances of AST breaks—between Block 1 and Block 2, it breaks If, FuncDef, and ClassDef; between Block 2 and Block 3, it breaks Attr, BinaryExpr, While, If, FuncDef, and ClassDef. In contrast, our AST-Aware approach results in only three breaks: between Block 1 and Block 2, it breaks ClassDef, and between Block 2 and Block 3, it breaks FuncDef and ClassDef.
To identify optimal partition boundaries, we use a dynamic programming (DP) algorithm:
@@
    The algorithm outputs the partition associated with dp[k_min, n], where $\texttt{k\_min}=\arg\min_{k}(dp[k,n])$, as the most optimal partition.
@@
# n: the length of the code file (number of tokens)
# m: the max number of segments; approximately n / max_len
@@
while q.nonempty() and q.left() < i - max_len:
q.pop_left() # pop indices before i - max_len
while q.nonempty() and dp[k - 1, q.right()] > dp[k - 1, i - 1]:
q.pop_right() # maintain monotonicity of values
@@
best_j = q.left() # guaranteed to have the smallest value
@@ 3.3 Pretraining with Span Corruption
The pretraining task of AST-T5 is based on span corruption, a well-established method for pretraining Transformer encoder-decoder models (Raffel et al., 2020). In this approach, 15% of the input tokens are randomly masked and replaced by unique “sentinel” tokens, distinct within each example. Each unique sentinel token is associated with a specific ID and added to the model’s vocabulary.
@@
3.4 Our AST-Aware Subtree Corruption
@@
def mask_subtree(t: ASTNode, m: int): # mask m tokens in subtree t
@@
m_remaining = m # distribute m tokens among children of t
@@
if child.size > theta: # a hyperparameter to control granularity
m_child = m * (child.size / t.size) # same mask ratio
mask_subtree(child, m_child) # mask recursively
@@
for child in ordered_children: # greedy allocation
@@
    Control the masking granularity via a tunable hyperparameter $\theta$ (named theta in Algorithm 2, Line 5).
The “mask quota” $m$ denotes the number of tokens to be masked in a subtree rooted at node $t$. The size of a subtree corresponds to the number of tokens it encompasses, derived from the cumulative sizes of its children. For larger subtrees that exceed the size threshold $\theta$, masking is applied recursively (Lines 5-8). Meanwhile, smaller subtrees undergo a weighted shuffle, and the quota $m$ is then apportioned among $t$’s children in a greedy fashion according to the shuffled order (Lines 11-15). The weights for shuffling are determined by a heuristic function on the size of each child, such that masking probabilities are distributed uniformly across leaf nodes. To create a subtree mask for an AST rooted at $t$ with a mask ratio $r$ (e.g., 15% or 25%), one can use $\texttt{mask\_subtree}(t,\lfloor|t|\cdot r\rfloor)$.
@@ 4 Experimental Setup
AST-T5 is pretrained on a mixed corpus consisting of code and natural language. Code is sourced from “GitHub repositories” dataset on Google BigQuery, which includes all code files from repositories with open-source licenses permitting redistribution. For NL, we use Wikipedia and OpenWebText, following Liu et al. (2019). Our corpus consists of 408 GB of code and 64 GB of text, smaller than the corpus used by CodeT5 (Wang et al., 2021) and CodeT5+ (Wang et al., 2023). Detailed statistics are provided in Section A.3.
Each code file is first parsed into its AST using the Tree-Sitter multi-language parser, and then tokenized with byte-level Byte-Pair Encoding (BPE) using a 64k BPE token vocabulary. Following AST-Aware Segmentation, these files are partitioned into chunks of 1,024 tokens. Our model is pretrained using the AST-Aware Subtree Corruption objective for 524 billion tokens (1,024 tokens per sequence, 1,024 sequences per batch, and 500k steps). For each training example, we apply AST-Aware Subtree Corruption of it is code, or apply Vanilla T5 Span Corruption of it is natural language. For code, the threshold, $\theta$, is uniformly sampled from 5 to 100. Pretraining uses PyTorch, Fairseq [^2] https://github.com/facebookresearch/fairseq and FlashAttention (Dao et al., 2022) and is conducted on 8 nodes, each with 8x NVIDIA A100 40GB GPUs. Further pretraining hyperparameters are detailed in Section A.4.
@@
We evaluate AST-T5 across three types of tasks: text-to-code generation, code-to-code transpilation, and code understanding (classification). Our evaluation encompasses tasks from the CodeXGLUE meta-benchmark (Lu et al., 2021) and also includes HumanEval (Chen et al., 2021a) and MBPP (Austin et al., 2021). Details about the benchmarks are shown in Section A.5.
@@ 5.1 Pretraining Procedure Analysis
+ Mask 25% (AST-T5) 12.8 22.9 23.8/16.1 68.9/72.3 98.6 65.8 47.7
+ Mask 50% 13.0 22.0 21.9/15.0 66.5/70.1 97.1 64.2 46.2
@@
The typical span corruption mask ratio in T5 is set at 15%. Increasing this ratio could potentially enhance the model’s generation capabilities, albeit potentially at the expense of understanding tasks. Essentially, a mask ratio of 100% would emulate a GPT-like, decoder-only Transformer. However, in our experiments (last two rows of Table 1), we observed that raising the mask ratio from 15% to 25% significantly improved generation capabilities without noticeably compromising performance in understanding tasks. Further analysis shows that increasing the masking ratio to 50% yields only a marginal improvement on HumanEval (from 12.8 to 13.0), while adversely impacting transpilation and understanding tasks. Thus, we settled on a 25% mask ratio for our AST-T5 model.
@@ 5.2 Main Results
AST-T5 (Ours) 12.8 22.9 23.8/16.1 68.9/72.3 98.6 65.8
Figure 3: Visualization of AST-T5’s performance on HumanEval compared to models exceeding 230M parameters and thus not detailed in Table 2. Each point on the scatter plot represents a model. The x-axis shows the parameter count in log-scale, while the y-axis shows the Pass@1 rate on HumanEval in log-scale. Model open-source status is color-coded: blue for open-source and red for proprietary.
Figure 4: Visualization of AST-T5’s performance on MBPP compared to other models. Each point on the scatter plot represents a model.
Table 2 shows AST-T5’s performance on downstream tasks compared with previously published results of similarly sized models, specifically those within the “Base” scale (110M to 230M parameters). Figure 3 and Figure 4 extends this comparison, comparing AST-T5 with larger models using the HumanEval benchmark and the MBPP benchmark, respectively. These results show that:
@@
While comparable in size, AST-T5 consistently outperforms similar-sized models such as CodeT5 (Wang et al., 2021) and CodeT5+ (Wang et al., 2023) in code generation, transpilation, and understanding. Notably, while CodeT5 and CodeT5+ are models at the Base scale, they were evaluated across different tasks. Our model, AST-T5, outperforms the best results of these two models across multiple benchmarks at the same time. Moreover, Figure 3 highlights AST-T5’s competitiveness against significantly larger models like GPT-J (Wang & Komatsuzaki, 2021) and LLaMa-7B (Touvron et al., 2023) on the HumanEval benchmark, underscoring our model’s parameter efficiency.
@@ A.2 More about AST-Aware Segmentation
Each element is only pushed into and popped out of the monotonic queue once, so the time complexity of the optimized algorithm is $O(nm)=O(n^{2}/\mathrm{max\_len})$, making the algorithm 1̃000x faster when $\mathrm{max\_len}=1024$. This allows the algorithm to segment each code file with 100k tokens in milliseconds.
A.3 Pretraining Corpora
Our pretraining corpora consists of two parts: code and natural language, coming from three sources:
  • GitHub (408 GB): The “GitHub repositories” public dataset available on Google BigQuery  [^3] https://console.cloud.google.com/marketplace/details/github/github-repos. For pretraining, we use all code files in Python (70 GB), C/C++ (195 GB), Java (105 GB), C# (38 GB) from each repo with an open-source license that explicitly permits redistribution.
  • Wikipedia (16 GB): A natural language corpus widely used for natural language pretraining.
  • OpenWebText (38 GB): A natural language corpus used by Liu et al. (2019) to train language models.
A.4 Pretraining Hyperparameters

### v2

@@ Abstract
Large language models (LLMs) have made significant advancements in code-related tasks, yet many LLMs treat code as simple sequences, neglecting its structured nature. We introduce AST-T5, a novel pretraining paradigm that leverages the Abstract Syntax Tree (AST) for enhanced code generation, transpilation, and understanding. Using dynamic programming, our AST-Aware Segmentation retains code structure, while our AST-Aware Span Corruption objective equips the model to reconstruct various code structures. Unlike other models, AST-T5 avoids complex program analyses or architectural changes, so it integrates seamlessly with any encoder-decoder Transformer. Evaluations show that AST-T5 consistently outperforms similar-sized LMs across various code-related tasks including HumanEval and MBPP. Structure-awareness makes AST-T5 particularly powerful in code-to-code tasks, surpassing CodeT5 by 2 points in exact match score for the Bugs2Fix task and by 3 points in exact match score for Java-C# Transpilation in CodeXGLUE. Our code and model are publicly available at https://annonymized .
PLACEHOLDER: Keywords
@@ 1 Introduction
Figure 1: Comparison of AST-Aware Subtree Corruption and Vanilla T5 using a Python factorial function. Both methods replace masked spans with sentinel tokens (special tokens added to the vocabulary, shown as [X], [Y], and [Z] in the figure), with output sequences containing the original masked tokens. Inputs and targets are shown in byte-pair encoding (BPE); for instance, “factorial” is encoded into “fact” and “orial”. Unlike Vanilla T5, which masks random spans without considering code structure, our approach specifically targets spans aligned with AST subtrees, like expressions and statements.
We have witnessed the transformative impact of large language models (LLMs) on various aspects of artificial intelligence in recent years (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023), especially in code generation and understanding (Feng et al., 2020; Wang et al., 2021; Rozière et al., 2023). By pretraining on massive code corpora such as the GitHub corpus, LLMs learns rich representations, thereby becoming powerful tools for various downstream applications such as text-to-code generation (Chen et al., 2021a; Austin et al., 2021; Iyer et al., 2018), code-to-code transpilation (Lu et al., 2021; Lachaux et al., 2020; Tufano et al., 2019), and code understanding (mapping code to classification labels) (Zhou et al., 2019; Svajlenko et al., 2014).
@@
In our experiments, AST-T5 consistently outperforms baselines in code generation, transpilation, and understanding tasks. Through controlled experiments, we empirically demonstrate that these advancements are attributed to our AST-aware pretraining techniques. Notably, AST-T5 not only outperforms similar-sized models like CodeT5 and CodeT5+ across various benchmarks but also remains competitive with, or occasionally even exceeds, the performance of much larger models using the HumanEval (Chen et al., 2021a) and the MBPP (Austin et al., 2021) benchmarks. Furthermore, the inherent AST-awareness of AST-T5 offers unique advantages in structure-sensitive tasks, such as code-to-code transpilation and Clone Detection, highlighting its effectiveness at capturing the structural nuances of code.
@@ 3.1 Parsing Code Into ASTs
Unlike traditional language models on code that handle code as simple sequences of subword tokens, AST-T5 leverages the Abstract Syntax Tree (AST) of code to gain semantic insights. For parsing purposes, we assume the provided code is syntactically valid—a reasonable assumption for tasks like code transpilation and understanding. Instead of the often computationally-intensive or infeasible methods of Control-Flow Analysis (CFA) or code execution (Guo et al., 2021; Tipirneni et al., 2023), our method only requires the code to be parsable. We use Tree-sitter, a multi-language parser, to construct the ASTs, where each subtree represents a consecutive span of subword tokens, and every leaf node represents an individual token.
3.2 AST-Aware Segmentation
@@
Segmentation in language model pretraining is a critical yet often overlooked aspect. Transformer LMs impose token limits on input sequences, making segmentation essential for fitting these inputs within the max_len constraint. A naive approach is Greedy Segmentation, where each chunk, except the last, contains exactly max_len tokens Figure 2 (Left). This strategy has been widely adopted in previous works, such as CodeT5 (Wang et al., 2021).
@@
AST-Aware Segmentation is our novel approach designed to preserve the AST structure of code during segmentation. Unlike Greedy Segmentation, which can indiscriminately fragment AST structures, our method strategically minimizes such disruptions. As illustrated in the example in Figure 2, Greedy Segmentation leads to nine instances of AST breaks—between Block 1 and Block 2, it breaks If, FuncDef, and ClassDef; between Block 2 and Block 3, it breaks Attr, BinaryExpr, While, If, FuncDef, and ClassDef. In contrast, our AST-Aware approach results in only three breaks: between Block 1 and Block 2, it breaks ClassDef, and between Block 2 and Block 3, it breaks FuncDef and ClassDef.
To identify optimal partition boundaries, we developed the following dynamic programming (DP)-based algorithm:
@@
    The algorithm outputs the partition associated with dp[k_min, n], where $\texttt{k\_min}=\arg\min_{k}(\texttt{dp}[k,n])$, as the most optimal partition.
@@
# n: the length of the code file
# (number of tokens)
# m: the max number of segments;
# approximately n / max_len
@@
while (q.nonempty() and
q.left() < i - max_len):
# pop indices before i - max_len
q.pop_left()
while (q.nonempty() and
dp[k-1, q.right()] > dp[k-1, i-1]):
# maintain monotonicity of values
q.pop_right()
@@
best_j = q.left()
# guaranteed to have the smallest value
@@ 3.3 Pretraining with Span Corruption
AST-T5’s pretraining is based on span corruption, a well-established method for pretraining transformer encoder-decoder models (Raffel et al., 2020). In this approach, 15% of the input tokens are randomly masked and replaced by unique “sentinel” tokens, distinct within each example. Each unique sentinel token is associated with a specific ID and added to the model’s vocabulary.
@@
3.4 AST-Aware Subtree Corruption
@@
def mask_subtree(t: ASTNode, m: int):
”””mask m tokens in subtree t”””
@@
m_remaining = m
# distribute m tokens among children of t
@@
# theta: a hyperparameter to control
# masking granularity
if child.size > theta:
# same mask ratio as the current subtree
m_child = m * (child.size / t.size)
mask_subtree(child, m_child) # recurse
@@
# greedy allocation of remaining mask quota
for child in ordered_children:
@@
    Control the masking granularity via a tunable hyperparameter $\theta$ (named theta in Algorithm 2, Line 9).
The “mask quota” $m$ denotes the number of tokens to be masked in a subtree rooted at node $t$. The size of a subtree corresponds to the number of tokens it encompasses, derived from the cumulative sizes of its children. For larger subtrees that exceed the size threshold $\theta$, masking is applied recursively (Lines 9-13). Meanwhile, smaller subtrees undergo a weighted shuffle, and the quota $m$ is then apportioned among $t$’s children in a greedy fashion according to the shuffled order (Lines 17-21). The weights for shuffling are determined by a heuristic function on the size of each child, such that masking probabilities are distributed uniformly across leaf nodes. To create a subtree mask for an AST rooted at $t$ with a mask ratio $r$ (e.g., 15% or 25%), one can use $\texttt{mask\_subtree}(t,\lfloor|t|\cdot r\rfloor)$.
@@ 4 Experimental Setup
AST-T5 is pretrained on a subset of The Stack Dedup corpus (Kocetkov et al., 2022), a near-deduplicated version of The Stack—a 3.1TB collection of permissively licensed source code from GitHub cutoff at April 2022, spanning 30 programming languages. For our experiments, AST-T5’s training involves Python, Java, C, C++, C#, and Markdown subsets, comprising a 588GB dataset with 93M code and natural language files.
Each file is first parsed into its AST using the Tree-Sitter multi-language parser, and then tokenized with byte-level Byte-Pair Encoding (BPE) using a 64k BPE token vocabulary. Following AST-Aware Segmentation, these files are partitioned into chunks of 1,024 tokens. Our model is pretrained using the AST-Aware Subtree Corruption objective for 524 billion tokens (1,024 tokens per sequence, 1,024 sequences per batch, and 500k steps). For each training example, we apply AST-Aware Subtree Corruption of it is code, or apply Vanilla T5 Span Corruption of it is natural language. For code, the threshold, $\theta$, is uniformly sampled from 5 to 100. Pretraining uses PyTorch, Fairseq [^2] https://github.com/facebookresearch/fairseq and FlashAttention (Dao et al., 2022) and is conducted on 8 nodes, each with 8x NVIDIA A100 40GB GPUs. Further pretraining hyperparameters are detailed in Section A.3.
@@
Table 1: Overview of our evaluation benchmarks about test set size, task type, and evaluation metric for each task. “Generation” tasks involve mapping natural language to code, “Transpilation” tasks involve translating code from one programming language to another, and “Understanding” tasks involve classifying code into categorical labels. For MBPP, we follow Nijkamp et al. (2023) and evaluate our model on the entire “sanitized” subset without few-shot prompts. For evaluation metrics, “Pass@1” indicates code execution on unit-tests provided in the benchmark using a single generated code per example, with reported pass rates. “EM” (Exact Match) evaluates textual equivalence without execution by comparing two canonicalized code pieces. “Acc” means accuracy in classification tasks. We omit “BLEU scores” because high BLEU values (>> 50) can still correspond to unexecutable or significantly flawed code (Lu et al., 2021), which is not useful in real-world applications. We also discuss evaluation results using the CodeBLEU (Ren et al., 2020) metric in Section A.5.
Size Type Metric
HumanEval 164 Generation Pass@1
MBPP 427 Generation Pass@1
Concode 2,000 Generation EM
Bugs2Fix 12,379 Transpilation EM
Java-C# 1,000 Transpilation EM
BigCloneBench 415,416 Understanding F1
Defect Detect 27,318 Understanding Acc
@@ 5 Evaluation Results
+ Mask 25% (AST-T5) 14.0 22.9 23.8/16.1 68.9/72.3 98.6 65.8 47.9
+ Mask 50% 14.3 22.0 21.9/15.0 66.5/70.1 97.1 64.2 46.4
5.1 Pretraining Procedure Analysis
@@
The typical span corruption mask ratio in T5 is set at 15%. Increasing this ratio could potentially enhance the model’s generation capabilities, albeit potentially at the expense of understanding tasks. Essentially, a mask ratio of 100% would emulate a GPT-like, decoder-only Transformer. However, in our experiments (last two rows of Table 2), we observed that raising the mask ratio from 15% to 25% significantly improved generation capabilities without noticeably compromising performance in understanding tasks. Further analysis shows that increasing the masking ratio to 50% yields only a marginal improvement on HumanEval (from 14.0 to 14.3), while adversely impacting transpilation and understanding tasks. Thus, we settled on a 25% mask ratio for our AST-T5 model.
@@ 5.2 Main Results
AST-T5 (Ours) 14.0 22.9 23.8/16.1 68.9/72.3 98.6 65.8
(a) HumanEval
(b) MBPP
Figure 3: Visualizations of AST-T5’s performance on HumanEval and MBPP compared to other models compared to models exceeding 230M parameters. Each point on each scatter plot represents a model. The x-axis shows the parameter count in log-scale, while the y-axis shows the Pass@1 rate on HumanEval or MBPP in log-scale. Model open-source status is color-coded: blue for open-source and red for proprietary.
Table 3 shows AST-T5’s performance on downstream tasks compared with previously published results of similarly sized models, specifically those within the “Base” scale (110M to 230M parameters). Figure 3(a) and Figure 3(b) extends this comparison, comparing AST-T5 with larger models using the HumanEval benchmark and the MBPP benchmark, respectively. These results show that:
@@
While comparable in size, AST-T5 consistently outperforms similar-sized models such as CodeT5 (Wang et al., 2021) and CodeT5+ (Wang et al., 2023) in code generation, transpilation, and understanding. Notably, while CodeT5 and CodeT5+ are models at the Base scale, they were evaluated across different tasks. Our model, AST-T5, outperforms the best results of these two models across multiple benchmarks at the same time. Moreover, Figure 3(a) highlights AST-T5’s competitiveness against significantly larger models like GPT-J (Wang & Komatsuzaki, 2021) and LLaMa-7B (Touvron et al., 2023) on the HumanEval benchmark, underscoring our model’s parameter efficiency. Similarly, Figure 3(b) demonstrates AST-T5’s advantages over LLaMa-7B and Codex-2.5B (Chen et al., 2021a) on the MBPP benchmark, showing the effectiveness of AST-T5.
@@ 6 Conclusion and Future Work
Broader Impact
In this paper, we introduce AST-T5, a language model aimed at automated generation, transpilation, and understanding of code. The advancement of LLMs in code generation raises concerns about automated code production’s security, privacy, and potential misuse. There is a risk that improved code generation capabilities could be exploited for malicious purposes, such as automating the creation of software vulnerabilities or facilitating the development of harmful software. Our research emphasizes the importance of responsible AI development and use, advocating for continuous monitoring, ethical guidelines, and safeguards to mitigate these risks.
@@ A.2 More about AST-Aware Segmentation
Each element is only pushed into and popped out of the monotonic queue once, so the time complexity of the optimized algorithm is $O(nm)=O(n^{2}/\mathrm{max\_len})$, making the algorithm $\sim$ 1000x faster when $\mathrm{max\_len}=1024$. This allows the algorithm to segment each code file with 100k tokens in milliseconds.
A.3 Pretraining Hyperparameters